# Kalman–Bucy Filtering with Randomized Sensing:

### v1

@@ V Application to Spatiotemporal Estimation
The continuous-time upper-bound framework developed in the previous section provides a general characterization of estimation uncertainty under stochastic sensing. We now consider the problem of spatiotemporal field estimation, where the process of interest is modeled as a GP evolving continuously in both space and time. To estimate the spatiotemporal field from noisy measurements collected at randomized sensor locations, we employ the  spatiotemporal Gaussian process Kalman filter (STGPKF) proposed in [9]. In this section, we demonstrate that the continuous-time upper-bound results derived earlier can be directly applied to the STGPKF, enabling a principled analysis of the fundamental limits of spatiotemporal estimation performance.
@@ V-B Spatiotemporal Gaussian Process Kalman Filter
Following previous work in [9], the temporal kernel $k_{T}$ admits an equivalent finite-dimensional SDE representation. At each $\bm{p}^{\delta}_{i}$, let $\bm{s}_{i}(t)\in\mathbb{R}^{n_{k}}$ denote the latent temporal state induced by the temporal kernel, where $n_{k}$ is the state dimension. Stacking yields $\bm{s}(t):=[\bm{s}_{1}(t)^{\top},\ldots,\bm{s}_{N_{g}^{\delta}}(t)^{\top}]^{\top}\in\mathbb{R}^{N_{g}^{\delta}n_{k}}$. The latent state evolves according to
@@
where $\bm{W}(t)=[W_{1}(t),\ldots,W_{N_{g}^{\delta}}(t)]^{\top}$ collects independent standard Wiener processes, $\bm{A}_{0}\in\mathbb{R}^{n_{k}\times n_{k}}$ and $\bm{B}_{0}\in\mathbb{R}^{n_{k}\times 1}$ are the state and diffusion matrices for a single grid point determined by the chosen temporal kernel, and $\bm{A}:=\bm{I}_{N_{g}^{\delta}}\otimes\bm{A}_{0}$, $\bm{B}:=\bm{I}_{N_{g}^{\delta}}\otimes\bm{B}_{0}$. The continuous-time process model in Eq. (37) is equivalent to Eq. (1) with the process noise covariance defined as $\bm{Q}_{c}=\bm{B}\bm{B}^{\top}$ according to [21]. For the Matérn-$\tfrac{1}{2}$ kernel ($n_{k}=1$), $\bm{A}_{0}$ and $\bm{B}_{0}$ reduce to scalars, corresponding exactly to the isotropic case discussed earlier, so the result of Theorem 1 applies directly.
@@
For the continuous-time process model and discrete measurement model defined in (37) and (LABEL:eq:measurement_model), we follow the approach in [9], which discretizes the continuous-time dynamics into an equivalent discrete-time state–space representation (see, e.g., [22]):
@@
As shown in [9], the resulting discrete-time KF provides the optimal state estimate, and Eq. (43)–(LABEL:eq:measurement_model) together define a standard linear discrete-time KF. Let $(\hat{\bm{s}}_{k|k},\,\bm{\Sigma}_{k|k})$ denote the posterior mean and covariance of the latent state at step $k$. Conditioned on the time-varying sensing pair $(\bm{H}_{k},\bm{V}_{k})$, the KF update is:
@@ V-C Estimation Performance
To evaluate the estimation performance of the STGPKF, we adopt clarity [9], which rescales the differential entropy of a continuous random variable onto the range $[0,1]$, with higher values indicating greater certainty. The differential entropy $h[\bm{Y}]$ of a continuous random variable $\bm{Y}\in\mathbb{R}^{n}$ is defined as:
@@ V-D Numerical Evaluation of STGPKF and Clarity
In this section, we validate the STGPKF and assess its estimation performance using numerical simulations. The spatiotemporal field, modeled as a GP with temporal and spatial Matérn-$\tfrac{1}{2}$ correlations, admits the following state-space representation [9]:
@@ V-E1 Discrete-Time Upper Bound for STGPKF
We first extend the discrete-time stochastic sensor selection framework of [8], which derives upper and lower bounds on the expected estimation covariance using the Riccati recursion and demonstrates their tightness in steady state. Although their formulation considers low-dimensional systems with single-sensor activation, its principles remain applicable. Here, we generalize this framework to our scenario, STGPKF, enabling analysis under multi-agent randomized sensing.
@@
However, explicitly evaluating this expectation is intractable, so we seek an upper bound. The following recursion, adapted from Gupta et al. [8], provides such a result.
@@
[8, Thm. 3] [Discrete-time Upper Bound] Let $N$ be the number of sensor deployments. If the $i$-th deployment is chosen at time step $k$ with probability $\pi^{k}_{i}$ independently, then the expected covariance is upper bounded by $\bm{\Delta}_{k+1}$, where $\bm{\Delta}_{k}$ is given by the recursion:
@@ V-E2 Continuous-time Upper Bound for STGPKF
To overcome the limitations of the discrete-time recursion, we now turn to a continuous-time formulation. Following the approach in [11], we express the continuous-time measurement model corresponding to the discrete formulation in Eq. (LABEL:eq:measurement_model) as
@@
To evaluate the effectiveness of our continuous-time bound, we compute the discrete-time expected covariance $\mathbb{E}[\bm{\Sigma}_{k}]$ by running the STGPKF over $30$ Monte Carlo trials with sensor locations randomized at each step, and compare it against (i) the discrete-time upper bound $\bm{\Delta}_{k}$ adapted from [8] and (ii) our analytically derived continuous-time upper bound $\bm{\Delta}(t)$. For visualization, we apply the linear operator $\mathcal{L}(\cdot)$ that maps a covariance matrix to the mean of its diagonal entries, and we report both the absolute error $\mathcal{L}(\bm{\Delta}_{k})-\mathcal{L}(\bm{\Delta}(t))$ and the relative error $\frac{\mathcal{L}(\bm{\Delta}_{k})-\mathcal{L}(\bm{\Delta}(t))}{\mathcal{L}(\bm{\Delta}_{k})}$ over time.
@@ V-F Analytical Characterization and Fundamental Limits
First, we derive a closed-form expression for the averaged information matrix $\bar{\bm{G}}$. Then, motivated by prior work [19], which investigates how system size affects estimation performance, we analyze the structure of the averaged information matrix $\bar{\bm{G}}$ to identify the factors fundamentally governing estimation accuracy. We aim to show that varying the number of grid points within the same compact domain does not fundamentally affect estimation performance; rather, it is dominated by intrinsic sensing parameters. As the grid resolution refines ($\delta\to 0$), the steady-state lower bound of the averaged expected clarity $\bar{q}_{\Delta^{\Pi}_{\infty}}$ becomes independent of discretization and, for a given temporal kernel, depends mainly on the number of agents, measurement noise, and sampling frequency.
@@ V-F2 Fundamental Performance Limits
Consider a continuous, symmetric, positive semi-definite kernel $k_{S}$ defined in Eq. (33) on a compact domain $D$. Assume the operator $T$ takes a function $f(x)$ as its argument and outputs a new function as:
$Tf(x):=\int_{D}k(x,y)f(y)\,dy,$ (69)
@@
$(T_{\delta}f)(p):=\frac{|D|}{N_{g}}\sum_{j=1}^{N_{g}}k_{S}(p,p_{j})f(p_{j}),N_{g}=\frac{|D|}{\delta^{d}}$ (72)
where $\{p_{j}\}_{j}^{N_{g}}\subset D$ are grid points and $k_{S}$ is a continuous, symmetric, positive-definite kernel defined in Eq. (33). Let $\{\nu_{i,\delta}\}_{i=1}^{N_{g}}$ and $\{\nu_{i}\}_{i=1}^{\infty}$ denote eigenvalues of $T_{\delta}$ and of continuous operator $T$ defined in Lemma 7, respectively. Then, as $\delta\to 0$ ($N_{g}\to\infty$),
@@
Consider the Matérn-$\tfrac{1}{2}$ kernel ($n_{k}=1$) with $\bm{C}=C_{0}\bm{I}$ and eigenvalues $\{\nu_{i}\}_{i=1}^{\infty}$ of the associated continuous kernel operator in Eq. (33) on a given spatial domain. Then, as the grid spacing $\delta\to 0$, the steady-state lower bound of the averaged expected clarity defined in Eq. (64) converges to a finite limit, i.e.,
$\lim_{\delta\to 0}\bar{q}_{\Delta^{\Pi}_{\infty}}=\frac{1}{1+C_{0}^{\,2}\sum_{i=1}^{\infty}\nu_{i}\gamma\bigl(\theta C_{0}^{2}\nu_{i}\bigr)}$ (74)
which depends solely on the sensing parameter
@@
Let the normalized spatial kernel matrix be
$\frac{\bm{K}_{gg}}{N_{g}}=\bm{U}\,\mathrm{diag}(\nu_{1,\delta},\ldots,\nu_{N_{g},\delta})\,\bm{U}^{\top}=\bm{U}\,\frac{\bm{\Lambda}_{K}}{N_{g}}\,\bm{U}^{\top}$ (76)
where $\bm{U}$ is orthogonal and $\nu_{i,\delta}$ are the eigenvalues of $\frac{\bm{K}_{gg}}{N_{g}}$. From Eq. (68),
$\bar{\bm{G}}=\theta C_{0}^{2}\bm{U}\frac{\Lambda_{K}}{N_{g}}\bm{U}^{T}=\bm{U}\mathrm{diag}\left(\theta C_{0}^{2}\nu_{1,\delta},\ldots,\theta C_{0}^{2}\nu_{N_{g},\delta}\right)\bm{U}^{\top}$ (77)
Define $\bm{\Lambda}_{\Delta}=\mathrm{diag}\big(\gamma(\theta C_{0}^{2}\nu_{1,\delta}),\ldots,\gamma(\theta C_{0}^{2}\nu_{N_{g},\delta})\big)$. According to Theorem 2,
$\bm{\Delta}_{\infty}=\bm{U}\mathrm{diag}\big(\gamma(\theta C_{0}^{2}\nu_{1,\delta}),\ldots,\gamma(\theta C_{0}^{2}\nu_{N_{g},\delta})\big)\bm{U}^{\top}.$ (78)
@@
$\frac{\bm{\Delta}^{\Pi}_{\infty}}{N_{g}}$ $=\frac{1}{N_{g}}C_{0}^{2}\bm{K}_{gg}^{1/2}\bm{\Delta}_{\infty}\bm{K}_{gg}^{1/2}$
@@
$=C_{0}^{2}\bm{U}(\frac{\bm{\Lambda}_{K}}{N_{g}}\bm{\Lambda}_{\Delta})\bm{U}^{\top}.$
Therefore, the eigenvalues of $\frac{\bm{\Delta}^{\Pi}_{\infty}}{N_{g}}$ are $\{\nu_{i,\delta}\gamma(\theta C_{0}^{2}\nu_{i,\delta})\}_{i}^{N_{g}}$. Hence,
$\frac{1}{N_{g}}\operatorname{tr}(\bm{\Delta}^{\Pi}_{\infty})=\sum_{i}^{N_{g}}\lambda_{i}(\frac{\bm{\Delta}^{\Pi}_{\infty}}{N_{g}})=\sum_{i}^{N_{g}}C_{0}^{2}\nu_{i,\delta}\gamma(\theta C_{0}^{2}\nu_{i,\delta})$
By Mercer’s theorem in Lemma 7 and its Nyström approximation in Lemma 8, $\{\nu_{i,\delta}\}_{i=1}^{N_{g}}$ are approximations of the eigenvalues $\{\nu_{i}\}_{i=1}^{\infty}$ of the continuous kernel operator, satisfying $\nu_{i,\delta}\to\nu_{i}$ as $\delta\to 0$ and $\sum_{i=1}^{\infty}\nu_{i}<\infty$. From Theorem 2, $\gamma(\cdot)$ is continuous and bounded, i.e.,
$0\leq\gamma(\lambda)\leq\gamma(0)=:C_{\gamma}<\infty,$ (79)
@@
$0\leq\nu_{i}\gamma(\theta C_{0}^{2}\nu_{i})\leq C_{\gamma}\nu_{i}.$ (80)
@@
$\lim_{\delta\to 0}\frac{1}{N_{g}}\,\operatorname{tr}\bigl(\bm{\Delta}^{\bm{\Pi}}_{\infty}\bigr)$ $=C_{0}^{\,2}\sum_{i=1}^{\infty}\nu_{i}\gamma\bigl(\theta C_{0}^{2}\nu_{i}\bigr)$
$\leq C_{0}^{\,2}\sum_{i=1}^{\infty}\nu_{i}C_{\gamma}$
converges to a finite constant.
Since $f_{q}=\frac{1}{1+x}$ is bounded and continuous on $0,\infty)$, the steady–state lower bound
@@
$=\frac{1}{1+C_{0}^{\,2}\sum_{i=1}^{\infty}\nu_{i}\gamma\bigl(\theta C_{0}^{2}\nu_{i}\bigr)}$
converges as $\delta\to 0$. When the environment (kernel parameters and domain) is fixed, the steady-state lower bound converges to a constant that depends solely on the sensing parameter $\theta$, independent of the grid resolution. ∎

### v2

@@ V Application to Spatiotemporal Estimation
The continuous-time upper-bound framework developed in the previous section provides a general characterization of estimation uncertainty under stochastic sensing. We now consider the problem of spatiotemporal field estimation, where the process of interest is modeled as a GP evolving continuously in both space and time. To estimate the spatiotemporal field from noisy measurements collected at randomized sensor locations, we employ the  spatiotemporal Gaussian process Kalman filter (STGPKF) proposed in [agrawal2024multi]. In this section, we demonstrate that the continuous-time upper-bound results derived earlier can be directly applied to the STGPKF, enabling a principled analysis of the fundamental limits of spatiotemporal estimation performance.
@@ V-B Spatiotemporal Gaussian Process Kalman Filter
Following previous work in [agrawal2024multi], the temporal kernel $k_{T}$ admits an equivalent finite-dimensional SDE representation. At each $\bm{p}^{\delta}_{i}$, let $\bm{s}_{i}(t)\in\mathbb{R}^{n_{k}}$ denote the latent temporal state induced by the temporal kernel, where $n_{k}$ is the state dimension. Stacking yields $\bm{s}(t):=[\bm{s}_{1}(t)^{\top},\ldots,\bm{s}_{N_{g}^{\delta}}(t)^{\top}]^{\top}\in\mathbb{R}^{N_{g}^{\delta}n_{k}}$. The latent state evolves according to
@@
where $\bm{W}(t)=[W_{1}(t),\ldots,W_{N_{g}^{\delta}}(t)]^{\top}$ collects independent standard Wiener processes, $\bm{A}_{0}\in\mathbb{R}^{n_{k}\times n_{k}}$ and $\bm{B}_{0}\in\mathbb{R}^{n_{k}\times 1}$ are the state and diffusion matrices for a single grid point determined by the chosen temporal kernel, and $\bm{A}:=\bm{I}_{N_{g}^{\delta}}\otimes\bm{A}_{0}$, $\bm{B}:=\bm{I}_{N_{g}^{\delta}}\otimes\bm{B}_{0}$. The continuous-time process model in Eq. (37) is equivalent to Eq. (1) with the process noise covariance defined as $\bm{Q}_{c}=\bm{B}\bm{B}^{\top}$ according to [sarkka2013gaussian]. For the Matérn-$\tfrac{1}{2}$ kernel ($n_{k}=1$), $\bm{A}_{0}$ and $\bm{B}_{0}$ reduce to scalars, corresponding exactly to the isotropic case discussed earlier, so the result of Theorem 1 applies directly.
@@
For the continuous-time process model and discrete measurement model defined in (37) and (LABEL:eq:measurement_model), we follow the approach in [agrawal2024multi], which discretizes the continuous-time dynamics into an equivalent discrete-time state–space representation (see, e.g., [wahlstrom2014discretizing]):
@@
As shown in [agrawal2024multi], the resulting discrete-time KF provides the optimal state estimate, and Eq. (43)–(LABEL:eq:measurement_model) together define a standard linear discrete-time KF. Let $(\hat{\bm{s}}_{k|k},\,\bm{\Sigma}_{k|k})$ denote the posterior mean and covariance of the latent state at step $k$. Conditioned on the time-varying sensing pair $(\bm{H}_{k},\bm{V}_{k})$, the KF update is:
@@ V-C Estimation Performance
To evaluate the estimation performance of the STGPKF, we adopt clarity [agrawal2024multi], which rescales the differential entropy of a continuous random variable onto the range $[0,1]$, with higher values indicating greater certainty. The differential entropy $h[\bm{Y}]$ of a continuous random variable $\bm{Y}\in\mathbb{R}^{n}$ is defined as:
@@ V-D Numerical Evaluation of STGPKF and Clarity
In this section, we validate the STGPKF and assess its estimation performance using numerical simulations. The spatiotemporal field, modeled as a GP with temporal and spatial Matérn-$\tfrac{1}{2}$ correlations, admits the following state-space representation [agrawal2024multi]:
@@ V-E1 Discrete-Time Upper Bound for STGPKF
We first extend the discrete-time stochastic sensor selection framework of [gupta2006stochastic], which derives upper and lower bounds on the expected estimation covariance using the Riccati recursion and demonstrates their tightness in steady state. Although their formulation considers low-dimensional systems with single-sensor activation, its principles remain applicable. Here, we generalize this framework to our scenario, STGPKF, enabling analysis under multi-agent randomized sensing.
@@
However, explicitly evaluating this expectation is intractable, so we seek an upper bound. The following recursion, adapted from Gupta et al. [gupta2006stochastic], provides such a result.
@@
[gupta2006stochastic, Thm. 3] [Discrete-time Upper Bound] Let $N$ be the number of sensor deployments. If the $i$-th deployment is chosen at time step $k$ with probability $\pi^{k}_{i}$ independently, then the expected covariance is upper bounded by $\bm{\Delta}_{k+1}$, where $\bm{\Delta}_{k}$ is given by the recursion:
@@ V-E2 Continuous-time Upper Bound for STGPKF
To overcome the limitations of the discrete-time recursion, we now turn to a continuous-time formulation. Following the approach in [lewis2017optimal], we express the continuous-time measurement model corresponding to the discrete formulation in Eq. (LABEL:eq:measurement_model) as
@@
To evaluate the effectiveness of our continuous-time bound, we compute the discrete-time expected covariance $\mathbb{E}[\bm{\Sigma}_{k}]$ by running the STGPKF over $30$ Monte Carlo trials with sensor locations randomized at each step, and compare it against (i) the discrete-time upper bound $\bm{\Delta}_{k}$ adapted from [gupta2006stochastic] and (ii) our analytically derived continuous-time upper bound $\bm{\Delta}(t)$. For visualization, we apply the linear operator $\mathcal{L}(\cdot)$ that maps a covariance matrix to the mean of its diagonal entries, and we report both the absolute error $\mathcal{L}(\bm{\Delta}_{k})-\mathcal{L}(\bm{\Delta}(t))$ and the relative error $\frac{\mathcal{L}(\bm{\Delta}_{k})-\mathcal{L}(\bm{\Delta}(t))}{\mathcal{L}(\bm{\Delta}_{k})}$ over time.
@@ V-F Analytical Characterization and Fundamental Limits
First, we derive a closed-form expression for the averaged information matrix $\bar{\bm{G}}$. Then, motivated by prior work [tzoumas2016sensor], which investigates how system size affects estimation performance, we analyze the structure of the averaged information matrix $\bar{\bm{G}}$ to identify the factors fundamentally governing estimation accuracy. We aim to show that varying the number of grid points within the same compact domain does not fundamentally affect estimation performance; rather, it is dominated by intrinsic sensing parameters. As the grid resolution refines ($\delta\to 0$), the steady-state lower bound of the averaged expected clarity $\bar{q}_{\Delta^{\Pi}_{\infty}}$ becomes independent of discretization and, for a given temporal kernel, depends mainly on the number of agents, measurement noise, and sampling frequency.
@@ V-F2 Fundamental Performance Limits
Consider a continuous, symmetric, positive semi-definite kernel $k_{S}$ defined in Eq. (33) on a compact domain $D$. Assume the operator $T$ takes a function $f(\cdot)$ as its argument and outputs a new function as:
$Tf(\bm{p}):=\int_{D}k_{S}(\bm{p},\bm{p}^{\prime})f(\bm{p}^{\prime})\,d\bm{p}^{\prime},$ (69)
@@
$(T_{\delta}f)(\bm{p}):=\frac{|D|}{N_{g}}\sum_{j=1}^{N_{g}}k_{S}(\bm{p},\bm{p}_{j})f(\bm{p}_{j}),N_{g}=\frac{|D|}{\delta^{d}}$ (72)
where $\{\bm{p}_{j}\}_{j}^{N_{g}}\subset D$ are grid points and $k_{S}$ is a continuous, symmetric, positive-definite kernel defined in Eq. (33). Let $\{\nu_{i,\delta}\}_{i=1}^{N_{g}}$ and $\{\nu_{i}\}_{i=1}^{\infty}$ denote eigenvalues of $T_{\delta}$ and of continuous operator $T$ defined in Lemma 7, respectively. Then, as $\delta\to 0$ ($N_{g}\to\infty$),
@@
Consider the Matérn-$\tfrac{1}{2}$ kernel ($n_{k}=1$) with $\bm{C}=C_{0}\bm{I}$ and eigenvalues $\{\nu_{i}\}_{i=1}^{\infty}$ of the associated continuous kernel operator in Eq. (69) on a given spatial domain. Define the function $\gamma:\mathbb{R}_{+}\to\mathbb{R}_{+}$ associated with the closed-form Riccati solution in Theorem 2 as
$\gamma(\lambda):=\frac{-q_{c}}{a-\sqrt{a^{2}+q_{c}\lambda}}.$ (74)
Then, as the grid spacing $\delta\to 0$, when the environment (kernel parameters and domain) is fixed, the steady-state lower bound of the averaged expected clarity defined in Eq. (64) converges to a finite limit, i.e.,
$\lim_{\delta\to 0}\bar{q}_{\Delta^{\Pi}_{\infty}}=\frac{1}{1+C_{0}^{\,2}\sum_{i=1}^{\infty}\frac{\nu_{i}}{|D|}\gamma\bigl(\theta C_{0}^{2}\frac{\nu_{i}}{|D|}\bigr)}$ (75)
which depends on the sensing parameter
@@
According to Lemma 8, the matrix form of the discrete kernel operator $T_{\delta}$ is $\frac{|D|\bm{K}_{gg}}{N_{g}}$. Since $\frac{|D|\bm{K}_{gg}}{N_{g}}\in\mathcal{S}_{++}^{N_{g}}$, it admits eigen-decomposition
$\frac{|D|\bm{K}_{gg}}{N_{g}}=\bm{U}\bm{\Lambda}\bm{U}^{\top},$
where $\bm{U}$ is an orthogonal matrix. For a discrete spatial domain with grid spacing $\delta$, we have the diagonal matrix $\bm{\Lambda}=\mathrm{diag}(\nu_{1,\delta},\ldots,\nu_{N_{g},\delta})$ as given in Lemma 8. Similarly,
$\bm{K}_{gg}=\bm{U}\bm{\Lambda}_{K}\bm{U}^{\top},$
where $\bm{\Lambda}_{K}:=\frac{N_{g}}{|D|}\bm{\Lambda}$. From Eq. (68), we know $\bar{\bm{G}}$ shares the same orthogonal matrix $\bm{U}$ as $\frac{\bm{K}_{gg}}{N_{g}}$, i.e.,
$\bar{\bm{G}}=\bm{U}\bm{\Lambda}_{G}\bm{U}^{\top},\,\bm{\Lambda}_{G}=\mathrm{diag}(\lambda_{1,\delta},\ldots,\lambda_{N_{g},\delta})$
where $\{\lambda_{i,\delta}\}_{i=1}^{N_{g}}$ are the eigenvalues of $\bar{\bm{G}}$ with each value $\lambda_{i,\delta}=\theta C_{0}^{2}\frac{\nu_{i,\delta}}{|D|}$. Then the eigenvalues $\gamma_{i,\delta}$ of $\bm{\Delta}_{\infty}$ satisfy $\gamma_{i,\delta}=\gamma(\lambda_{i,\delta})=\gamma(\theta C_{0}^{2}\frac{\nu_{i,\delta}}{|D|})$. According to Theorem 2, $\bm{\Delta}_{\infty}$ also shares the same orthogonal matrix $\bm{U}$ as $\bar{\bm{G}}$, i.e.,
$\bm{\Delta}_{\infty}$ $=\bm{U}\bm{\Lambda}_{\Delta}\bm{U}^{\top},\,\bm{\Lambda}_{\Delta}=\mathrm{diag}\big(\gamma_{1,\delta},\ldots,\gamma_{N_{g},\delta}\big).$
@@
$\frac{\bm{\Delta}^{\Pi}_{\infty}}{N_{g}}$ $=\frac{1}{N_{g}}\mathcal{T}(\bm{\Delta}_{\infty})$
$=\frac{1}{N_{g}}C_{0}^{2}\bm{K}_{gg}^{1/2}\bm{\Delta}_{\infty}\bm{K}_{gg}^{1/2}$
@@
$=C_{0}^{2}\bm{U}(\frac{\bm{\Lambda}_{K}}{N_{g}}\bm{\Lambda}_{\Delta})\bm{U}^{\top}$
$=C_{0}^{2}\bm{U}(\frac{\bm{\Lambda}}{|D|}\bm{\Lambda}_{\Delta})\bm{U}^{\top}.$
Therefore, the eigenvalues of $\frac{\bm{\Delta}^{\Pi}_{\infty}}{N_{g}}$ are $\{\frac{\nu_{i,\delta}}{|D|}\gamma_{i,\delta}\}_{i=1}^{N_{g}}$. Hence,
$\frac{1}{N_{g}}\operatorname{tr}(\bm{\Delta}^{\Pi}_{\infty})=\sum_{i=1}^{N_{g}}C_{0}^{2}\frac{\nu_{i,\delta}}{|D|}\gamma_{i,\delta}=C_{0}^{2}\sum_{i=1}^{N_{g}}\frac{\nu_{i,\delta}}{|D|}\gamma(\theta C_{0}^{2}\frac{\nu_{i,\delta}}{|D|})$
By Mercer’s theorem in Lemma 7 and its Nyström approximation in Lemma 8, $\{\nu_{i,\delta}\}_{i=1}^{N_{g}}$ are approximations of the eigenvalues $\{\nu_{i}\}_{i=1}^{\infty}$ of the continuous kernel operator, satisfying $\nu_{i,\delta}\to\nu_{i}$ as $\delta\to 0$ and $\sum_{i=1}^{\infty}\nu_{i}<\infty$. In addition, $\gamma(\cdot)$ is continuous and bounded, i.e.,
$0\leq\gamma(\lambda)\leq\gamma(0)=:C_{\gamma}<\infty,$
@@
$0\leq\frac{\nu_{i}}{|D|}\gamma(\theta C_{0}^{2}\frac{\nu_{i}}{|D|})\leq\frac{\nu_{i}}{|D|}C_{\gamma}.$
@@
$\lim_{\delta\to 0}\frac{1}{N_{g}}\,\operatorname{tr}\bigl(\bm{\Delta}^{\bm{\Pi}}_{\infty}\bigr)$ $=C_{0}^{\,2}\sum_{i=1}^{\infty}\frac{\nu_{i}}{|D|}\gamma\bigl(\theta C_{0}^{2}\frac{\nu_{i}}{|D|}\bigr)$
$\leq C_{0}^{\,2}\sum_{i=1}^{\infty}\frac{\nu_{i}}{|D|}C_{\gamma}$
converges to a finite constant. Since $f_{q}(x)=\frac{1}{1+x}$ is bounded and continuous on $0,\infty)$, the steady–state lower bound
@@
$=\frac{1}{1+C_{0}^{\,2}\sum_{i=1}^{\infty}\frac{\nu_{i}}{|D|}\gamma\bigl(\theta C_{0}^{2}\frac{\nu_{i}}{|D|}\bigr)}$
also converges as $\delta\to 0$. When the environment (kernel parameters and domain) is fixed, the steady-state lower bound converges to a constant that depends solely on the sensing parameter $\theta$, independent of the grid resolution. ∎